n identical cells , each of emf E and internal resistance r , are joined in series to form  a closed circuit. one of the cell A is joined with reversed polarity the potential difference across each cell except A  is :-  plzz give proper solution with detail explanation plzzzzzzzzzzzzz

Grade:11

2 Answers

Arun
6 years ago
I = (ne- 2e)/nr
I = (n-2)e/nr
As we know
V = e - Ir
V = e - (n -2)e/n
V = (ne - ne +2e)/n
V = 2e/n
Hope you got the answer
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to the question.
 
The potential drop across circuit, V = (n – 1)E – E = (n – 2)E
Current flowing in the circuit, I = (n – 2)E/nr
Hence potential drop across each cell (except A), V’ = E – Ir
                                                                             = E – (n – 2)E/n
                                                                             = 2E/n
Hope it helps.
Thanks and regards,
Kushagra